Find the magnitude and direction of -A+B where A=(23.0,59.0) , B = ( 90.0, -150.0 )


1
Expert's answer
2019-09-12T10:40:33-0400

The resulting vector "\bold{C} = \bold{B} - \bold{A} = (67.0, -209.0)" . The magnitude of this vector is "\| \bold{C} \| = \sqrt{67^2 + 209^2} \approx 219.48". The angle between "\bold{C}" and x-axis is "\theta = \arctan \frac{-209.0}{67.0} \approx -72.2 ^\circ" ("72.2^\circ" clockwise from the East direction).